OpenStudy (anonymous):

Find the roots of the polynomial given below. 3x2 - x - 5

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x^2 - x - 5 = 0}\)

OpenStudy (agent47):

3x^2-x-5=0 D=1+15=16

OpenStudy (agent47):

x1=(1-sqrt(16))/6

OpenStudy (agent47):

x2=(1+sqrt(16))/6

OpenStudy (agent47):

x1=-1/2 x2=5/6

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {x = {-b \pm\sqrt{b^2 - 4ac} \over 2a}}}\)